# Understanding Flatness in Generative Models: Its Role and Benefits

### v1

@@ 4.1 Experiments Settings
Experimental details. We trained the model for 200K steps on the CIFAR-10 and LSUN Tower datasets, and 120K steps on the FFHQ dataset. We employed the Adam optimizer for all experiments with a learning rate of $1e^{-}{4}$, following  [7, 30]. For +IP, we used an input perturbation strength of 0.1 following [32], and for SWA, we set the cycle length $c$ to 100 and start averaging from 180K steps for CIFAR-10 and LSUN Tower datasets. Finally, for +SAM, we tuned the $\rho\in[1e^{-1},1e^{-2},1e^{-3}]$ parameters.
@@ 4.2 Generative Performance
The quantitative and qualitative results presented in Table. 2 and Figure. 4 demonstrate the generative performance of baselines. From Table. 2, we observe that applying EMA consistently improves FID scores across all datasets and respaced timesteps, demonstrating its effectiveness in stabilizing model updates and enhancing sample quality. SWA also provides notable improvements but does not outperform EMA in most cases. The introduction of SAM, however, leads to the lowest FID scores, particularly in the +SAM+SWA setting, suggesting that explicitly controlling flatness results in a more robust and generalizable generative model. The trend is particularly evident in the 100-step respacing setting, where SAM-based models consistently outperform baselines, indicating that flat minima contribute to better sample quality, especially under reduced sampling steps. Figure. 4 further supports these findings by providing visual comparisons of generated images.
@@ 4.3 Flatness Measurement
In Table. 3, we compare LPF flatness measures [2], where lower values indicate a flatter loss landscape. We observe that ADM already possesses a certain level of flatness. Interestingly, we find that ensemble-like averaging, such as +SWA and +EMA, fail to induce additional flatness. In contrast, +SAM finds a flatter loss surface by explicitly perturbing model weights, as evidenced by SAM-applied models consistently achieving the lowest LPF values and significantly superior performance to other baselines. Also, we plot another measurement to support these findings further. Figure. 5 shows how the loss value increases as the model perturbation is imposed for CIFAR-10. +IP, +EMA, and +SWA show no significant differences, coincide with the result of LPF, where these fail to find additional flatness. In this setting, +SAM shows a flatter behavior than other algorithms. It shows that diffusion models already show flatter loss landscapes, and ensemble-like averaging, such as +SWA and EMA, shows less impact. Moreover, finding flat minima explicitly leads to the flatness in diffusion models. We provide plots for other algorithms, including +IP+EMA, +IP+SWA, +SAM+EMA, +SAM+SWA, in the Appendix.
@@ 4.4 Flat Diffusion Models are Robust
Exposure bias. In Figure. 6, we compare the L2 norm of the predicted noise in diffusion models. A model that closely aligns with the training trajectory is expected to exhibit lower exposure bias, thereby mitigating distribution shifts over iterative sampling steps. We observe that SAM-applied models consistently maintain L2 norms closer to the training curve, indicating a lower degree of exposure bias. This result aligns well with their superior FID performance reported in Table. 2, suggesting that flat diffusion models generalize better across sampling timesteps and effectively combat error accumulation.
Model quantazation error. Table. 4 presents the FID scores of quantization across baselines, when reducing precision from 32-bit to 8-bit. Here, we quantize model parameters to 8-bit values using scaling and rounding, without any additional fine-tuning or retraining. Thanks to +SAM that aims to the robustness of model perturbation, flatter diffusion models results in surprisingly better robustness under quantization. These findings also support that +SAM is robust to model quantization.
@@ 5 Related works
Diffusion models. Diffusion probabilistic models (DPMs) [12, 7, 38, 39, 41] have been recently emerged as powerful generative frameworks, achieving state-of-the-art performance in image [35, 33] and video [15, 3] synthesis. Despite extensive research on improving diffusion models through architecture modifications [29, 14, 33], enhanced training strategies [43], and novel noise scheduling techniques [41, 42], relatively little attention has been given to understanding their loss landscape properties.
Flat minima in various tasks. Flat minima have garnered significant attention in classification and domain generalization tasks with the view of loss landscape  [13, 9, 21] and objective function [8, 17]. Other tasks also adopt the flatness for their performance. In federated learning, for example, where client devices collect and train on heterogeneous datasets, ensuring robust generalization across diverse data distributions is a core challenge. Recent works such as [34, 4, 20] have employed flat minima strategies to mitigate this data heterogeneity issue by obtaining the model robust to distribution shift. Some findings encouraging flat reward in RL leads to learning of robust policies [19, 18].
Flat minima on diffusion models. Recent empirical findings suggest that diffusion models inherently exhibit flatter loss landscape [45]. This observation aligns with the broader literature on the benefits of flat minima, particularly in supervised learning, where flatter loss surfaces are known to enhance generalization and improve robustness to perturbations. However, the implications of flat minima in the context of diffusion models remain largely unexplored. While some prior works have attempted to leverage smooth function properties to improve generative modeling [1, 6, 32], they primarily focus on enforcing Lipschitz continuity or reducing gradient penalties rather than explicitly studying the role of flatness in diffusion models. To the best of our knowledge, no prior research has systematically investigated how flat minima influence diffusion models, how they impact generative performance, or whether they can be intentionally optimized to enhance diffusion model robustness. This gap in understanding motivates us, where we aim to establish a theoretical and empirical foundation for analyzing the effect of flat minima on diffusion models.
32-bit $\rightarrow$ 8-bit 20 steps 100 steps
@@
[…]
Table 4: FID performance on CIFAR-10 when model parameters are quantized to 8-bit. We apply direct quantization by clipping and rounding 32-bit parameters without additional fine-tuning. The 32-bit performance results are identical to those in Table. 2.
6 Conclusion
We presented an in-depth investigation of the role and benefits of flat loss surfaces in generative models, which have remained unknown despite the great success of deep generative models. Our study explored this unknown by analyzing how flatness influences generative modeling and demonstrating its possible impact on robustness. First, our theoretical analysis reveals that the loss flatness of the generative models enhances the robustness against the disruptions of the target data distributions or the perturbations of model parameters. These insights naturally link to the reduced exposure bias problem and the model quantization error. Based on the evaluations on CIFAR-10, LSUN Tower, and FFHQ, we demonstrated that 1) SAM is suitable to seek flatter minima of diffusion models rather than SWA and EMA, 2) flat minima reduce exposure bias with a minimal gap between testing and training, and 3) a flat model keeps its generative performance even with a strong model quantization; coinciding with our theory.
Supplementary material
Appendix A Mathematical claims and proofs
For the main claims, we follow $\mathcal{L}(\mathbf{x};\boldsymbol{\theta},t,p_{t}):=||s_{\boldsymbol{\theta}}(\mathbf{x},t)-\nabla_{\mathbf{x}}\log{p_{t}(\mathbf{x})}||_{2}^{2}$, while dropping the timestep $t$ without loss of generality. Our mathematical claims are valid for all timesteps.
@@ Appendix A Mathematical claims and proofs
where $l^{*}:=\mathcal{L}(\mathbf{x};\boldsymbol{\theta}^{*},p)$ and $\Delta\in\mathbb{R}^{+}$.333$\forall$ means ‘for all,’ $\exists$ means ‘there exists,’ and $\mathbb{R}^{+}$ indicates the set of positive real numbers
@@
Theorem 1. (A perturbed distribution) For a given prior distribution of $p(\mathbf{x})$ and the $\delta$-perturbed minimum, i.e., $\boldsymbol{\theta}+\delta$, the following $\hat{p}(\mathbf{x})$ satisfies the $\mathcal{L}(\mathbf{x};\boldsymbol{\theta}+\delta,p)=\mathcal{L}(\mathbf{x};\boldsymbol{\theta},\hat{p})$:
@@
[…]
@@
$s_{\boldsymbol{\theta}}(\mathbf{x},t):=\frac{1}{m}\boldsymbol{\theta}\sigma(\mathbf{W}^{T}\mathbf{x}+\mathbf{U}^{T}\mathbf{e}_{t})$ (22)
@@
$\mathcal{L}(\mathbf{x};\boldsymbol{\theta},p):=||s_{\boldsymbol{\theta}}(\mathbf{x})-\nabla_{\mathbf{x}}\log{p(\mathbf{x})}||_{2}^{2},$ (23)
@@
$\mathcal{L}(\mathbf{x};\boldsymbol{\theta}+\delta,p):=||s_{\boldsymbol{\theta}+\delta}(\mathbf{x})-\nabla_{\mathbf{x}}\log{p(\mathbf{x})}||_{2}^{2}.$ (24)
@@
$=\frac{1}{m}(\boldsymbol{\theta}+\delta)\sigma(\mathbf{W}^{T}\mathbf{x}+\mathbf{U}^{T}\mathbf{e})-\nabla_{\mathbf{x}}\log{p(\mathbf{x})}$ (26)
$=\frac{1}{m}\boldsymbol{\theta}\sigma(\mathbf{W}^{T}\mathbf{x}+\mathbf{U}^{T}\mathbf{e})+\frac{1}{m}\delta\sigma(\mathbf{W}^{T}\mathbf{x}+\mathbf{U}^{T}\mathbf{e})-\nabla_{\mathbf{x}}\log{p(\mathbf{x})}$ (27)
@@
$\frac{1}{m}\delta(\mathbf{W}^{T}\mathbf{x}+\mathbf{U}^{T}\mathbf{e})-\nabla_{\mathbf{x}}\log{p(\mathbf{x})}$ (28)
@@
$\frac{\partial I(\mathbf{x})}{\partial\mathbf{x}}=\frac{1}{m}\delta(\mathbf{W}^{T}\mathbf{x}+\mathbf{U}^{T}\mathbf{e})$ (29)
@@
$I(\mathbf{x})=\frac{1}{2}\mathbf{x}^{T}(\delta\mathbf{W}^{T})\mathbf{x}+\mathbf{x}^{T}\delta(\mathbf{U}^{T}\mathbf{e})+C,$ (30)
with the assumption $\delta\mathbf{W}^{T}$ is symmetric and where $C$ is a constant real number.
$\frac{1}{m}\delta(\mathbf{W}^{T}\mathbf{x}+\mathbf{U}^{T}\mathbf{e})-\nabla_{\mathbf{x}}\log{p(\mathbf{x})}$ (31)
@@
$=\exp{\Big{\{}-\frac{1}{2m}\mathbf{x}^{T}(\delta\mathbf{W}^{T})\mathbf{x}-\frac{1}{m}\mathbf{x}^{T}\delta(\mathbf{U}^{T}\mathbf{e})-C^{*})\Big{\}}}p(\mathbf{x})$ (36)
Corollary 1. (Diffusion version of Theorem 1) For a given prior Gaussian distribution of noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ and the $\delta$-perturbed minimum, i.e., $\boldsymbol{\theta}+\delta$, the following $\hat{\epsilon}$ satisfies the $\mathcal{L}(\mathbf{x};\boldsymbol{\theta}+\delta,p)=\mathcal{L}(\mathbf{x};\boldsymbol{\theta},\hat{p})$:
@@
  • $p(\mathbf{x})$: the true data distribution that is known in the training process.
  • $\hat{p}(\mathbf{x})$: the perturbed distribution that caused by the $\mathbf{\epsilon}$ model perturbation in Eq. (35).
@@
$\hat{\epsilon}_{t}(\mathbf{x})$ $:=e^{-I(\mathbf{x},\delta)}\epsilon_{t}$ (38)
$=\exp\bigg{(}-\frac{1}{2m}\mathbf{x}^{T}\delta_{w}\mathbf{x}-\frac{1}{m}\mathbf{x}^{T}\delta_{u}-C\bigg{)}\cdot\frac{1}{\sqrt{(2\pi)^{d}}}\exp\bigg{(}-\frac{1}{2}\mathbf{x}^{T}\mathbf{x}\bigg{)},$ (39)
[…]
$=\frac{1}{\sqrt{(2\pi)^{d}}}\exp\bigg{(}-\frac{1}{2}\mathbf{x}^{T}(\mathbf{I}+\frac{1}{m}\mathbf{\delta}_{w})\mathbf{x}-\frac{1}{m}\mathbf{x}^{T}\delta_{u}-C\bigg{)}$ (40)
$=\frac{1}{C^{\prime}}\exp\bigg{(}-\frac{1}{2}\bigg{\{}\mathbf{x}^{T}(\mathbf{I}+\frac{\delta_{w}}{m})\mathbf{x}+\frac{2}{m}\mathbf{x}^{T}\delta_{u}\bigg{\}}\bigg{)},C^{\prime}:=\frac{1}{(\sqrt{2\pi})^{d}}\exp(-C)$ (41)
$=\frac{1}{\sqrt{(2\pi)^{d}|\Sigma_{\delta}|}}\exp\bigg{(}-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu}_{\delta})^{T}{\Sigma_{\delta}}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{\delta})\bigg{)},\text{ where }{\Sigma_{\delta}}:=\bigg{(}\mathbf{I}+\frac{\delta_{w}}{m}\bigg{)}^{-1},\text{ and }\boldsymbol{\mu}_{\delta}:=\frac{1}{m}\Sigma_{\delta}\delta_{u}$ (42)
Because the $\hat{\epsilon}(\mathbf{x})$ is also the Gaussian distribution, we present the KL Divergence between $p(\mathbf{x})$ and $p^{*}_{\epsilon}(\mathbf{x})$ as follows:
$D_{KL}(\hat{\epsilon}\|\epsilon)$ $=\frac{1}{2}\Bigg{[}\log\frac{|\Sigma_{\delta}|}{|\Sigma|}-d+tr({\Sigma_{\delta}}^{-1}{\Sigma}^{-1})+{\boldsymbol{\mu}_{\delta}}^{T}{\Sigma_{\delta}}^{-1}\boldsymbol{\mu}_{\delta}\Bigg{]}$ (43)
$=\frac{1}{2}\Bigg{[}\log|\Sigma_{\delta}|-d+tr({\Sigma_{\delta}}^{-1})+{\boldsymbol{\mu}_{\delta}}^{T}{\Sigma_{\delta}}^{-1}\boldsymbol{\mu}_{\delta}\Bigg{]},\because\Sigma=\mathbf{I}$ (44)
@@
$\mathcal{E}$ $\leq\operatorname*{arg\,max}_{\|\delta\|_{2}\leq\Delta}\frac{1}{2}\Bigg{[}\log|\Sigma_{\delta}|-d+tr(\Sigma_{\delta}^{-1})+{\boldsymbol{\mu}_{\delta}}^{T}{\Sigma_{\delta}}^{-1}\boldsymbol{\mu}_{\delta}\Bigg{]}$ (49)
$\leq\frac{1}{2}\Bigg{[}\sum_{i}^{d}\log\frac{m}{m+\sigma_{i}}+\sum_{i}^{d}\frac{\sigma_{i}}{m}+\frac{\sigma_{d}}{m^{2}}\|\mathbf{U}^{T}\mathbf{e}\|_{2}^{2}\Delta^{2}\Bigg{]},$ (50)
@@
$\mathcal{E}$ $\leq\operatorname*{arg\,max}_{\|\delta\|_{2}\leq\Delta}\frac{1}{2}\Bigg{[}\log|\Sigma_{\delta}|-d+tr(\Sigma_{\delta}^{-1})+{\boldsymbol{\mu}_{\delta}}^{T}{\Sigma_{\delta}}^{-1}\boldsymbol{\mu}_{\delta}\Bigg{]}$ (51)
$\leq\frac{1}{2}\Bigg{[}\sum_{i}^{d}\log\frac{m}{m+\sigma_{i}}-d+\sum_{i}^{d}\bigg{(}1+\frac{\sigma_{i}}{m}\bigg{)}+\frac{\sigma_{d}}{m^{2}}\|\mathbf{U}^{T}\mathbf{e}\|_{2}^{2}\Delta^{2}$ (52)
$\leq\frac{1}{2}\Bigg{[}\sum_{i}^{d}\log\frac{m}{m+\sigma_{i}}+\sum_{i}^{d}\frac{\sigma_{i}}{m}+\frac{\sigma_{d}}{m^{2}}\|\mathbf{U}^{T}\mathbf{e}\|_{2}^{2}\Delta^{2}\Bigg{]},$ (53)

### v2

@@ 4 Experiments
Figure 4: Generation results of (a) full precision (32-bit) and (b) 8-bit quantized models. We use respaced timesteps $T^{\prime}=20$ for sampling. For the 8-bit case, we selectively compare +EMA, +SAM, and their combinations, where the relatively flatter minima have been found.
@@ 4.1 Experiments Settings
Experimental details. We trained the model for 200K steps on the CIFAR-10, LSUN Tower, and FFHQ datasets. We employed the Adam optimizer for all experiments with a learning rate of $1e^{-4}$, following  [7, 30]. For +IP, we used an input perturbation strength of 0.1 following [32], and we set the cycle length $c$ to 100 and start averaging from 180K steps for SWA. Finally, we tuned the $\rho\in[1e^{-1},1e^{-2},1e^{-3}]$ parameters for +SAM.
@@ 4.2 Generative Performance
The quantitative and qualitative results presented in Tabs. 2 and 4 demonstrate the generative performance of baselines. From Tab. 2, we observe that applying EMA consistently improves FID scores across all datasets and respaced timesteps, demonstrating its effectiveness in stabilizing model updates and enhancing sample quality. SWA also provides notable improvements but does not outperform EMA in most cases. [^3] Because of limited space, we include the discussion and experimental results of the post-hoc EMA (pEMA) [8] in Appendix B.3. While EMA and SWA can always be applied as auxiliary options, both standalone SAM and +SAM+SWA achieve comparable or better FID scores, especially under 20 timesteps, where exposure bias worsens. This suggests that explicit flatness control improves generalization, in terms of enhancing sample quality and robustness. For IP, the FID score is comparable to that of SAM at 100 steps, since the perturbation in the data distribution translates into parameter perturbations (Theorem 1). However, IP performs poorly at 20 steps, as it lacks a principled noise direction and tends to converge to suboptimal minima. Figure 4 further supports these findings by providing visual comparisons of generated images  [^4] We include additional sampling images in Appendix B.2.
@@ 4.3 Flatness Measurement
In Tab. 3, we compare LPF flatness measures [2], where lower values indicate a flatter loss landscape. We observe that ADM already possesses a certain level of flatness. Interestingly, we find that ensemble-like averaging, such as +SWA and +EMA, fail to induce additional flatness. In contrast, +SAM finds a flatter loss surface by explicitly perturbing model weights, as evidenced by SAM-applied models consistently achieving the lowest LPF values and significantly superior performance to other baselines. Also, we plot another measurement to support these findings further. Figure 5 shows how the loss value increases as the model perturbation is imposed for CIFAR-10.
+IP, +EMA, and +SWA show no significant differences, coincide with the result of LPF, where these fail to find additional flatness. In this setting, +SAM shows a flatter behavior than other algorithms. It shows that diffusion models already show flatter loss landscapes, and ensemble-like averaging, such as +SWA and EMA, shows less impact. Moreover, finding flat minima explicitly leads to the flatness in diffusion models. We provide plots for other algorithms, including +IP+EMA, +IP+SWA, +SAM+EMA, +SAM+SWA, in the Appendix B.1.
@@ 4.4 Flat Diffusion Models are Robust
Exposure bias. In Fig. 6, we compare the L2 norm of the predicted noise in diffusion models. A model that closely aligns with the training trajectory is expected to exhibit lower exposure bias, thereby mitigating distribution shifts over iterative sampling steps. We observe that SAM-applied models consistently maintain L2 norms closer to the training curve, indicating a lower degree of exposure bias. This result aligns well with their superior FID performance reported in  Tab. 2, suggesting that flat diffusion models generalize better across sampling timesteps and effectively combat error accumulation.
32-bit $\rightarrow$ 8-bit 20 steps 100 steps
ADM 34.47 $\overset{\text{{\tiny{\color[rgb]{1,0,0}+13.65}}}}{\longrightarrow}$ 48.02 8.80 $\overset{\text{{\tiny{\color[rgb]{1,0,0}+3.98}}}}{\longrightarrow}$ 12.78
+EMA 10.63 $\overset{\text{{\tiny{\color[rgb]{1,0,0}+10.02}}}}{\longrightarrow}$ 20.65 4.06 $\overset{\text{{\tiny{\color[rgb]{1,0,0}+3.3}}}}{\longrightarrow}$ 7.36
+SAM 9.01 -$\overset{\text{{\tiny{\color[rgb]{0,0.6,0}-0.07}}}}{\longrightarrow}$ 8.94 3.83 $\overset{\text{{\tiny{\color[rgb]{1,0,0}+0.19}}}}{\longrightarrow}$ 4.02
+SAM+EMA 7.00 $\overset{\text{{\tiny{\color[rgb]{1,0,0}+0.2}}}}{\longrightarrow}$ 7.20 3.18 -$\overset{\text{{\tiny{\color[rgb]{0,0.6,0}-0.06}}}}{\longrightarrow}$ 3.12
Table 4: FID performance on CIFAR-10 when model parameters are quantized to 8-bit. We apply direct quantization by clipping and rounding 32-bit parameters without additional fine-tuning. The 32-bit performance results are identical to those in  Table 2.
Model quantization error. Tab. 4 presents the FID scores of quantization across baselines, when reducing precision from 32-bit to 8-bit. Here, we quantize model parameters to 8-bit values using scaling and rounding, without any additional fine-tuning or retraining. Thanks to +SAM that aims to the robustness of model perturbation, flatter diffusion models results in surprisingly better robustness under quantization. These findings also support that +SAM is robust to model quantization.
@@ 5 Related works
Diffusion models. Diffusion probabilistic models (DPMs) [13, 7, 39, 40, 42] have recently emerged as powerful generative frameworks, achieving state-of-the-art performance in image [36, 33] and video [16, 3] synthesis. Despite extensive research on improving diffusion models through architecture modifications [29, 15, 33], enhanced training strategies [44], and novel noise scheduling techniques [42, 43], relatively little attention has been given to understanding their loss landscape properties.
Flat minima in various tasks. Flat minima have garnered significant attention in classification and domain generalization tasks with the view of loss landscape  [14, 10, 20] and objective function [9, 18]. Other tasks also adopt the flatness for their performance. In federated learning, for example, where client devices collect and train on heterogeneous datasets, ensuring robust generalization across diverse data distributions is a core challenge. Recent works have employed flat minima strategies to mitigate this data heterogeneity issue by obtaining the model robust to distribution shift [34, 4, 19] and enhance the robustness of compressed neural networks [28, 48].
Flat minima on diffusion models. Recent findings suggest that diffusion models inherently exhibit a flatter loss landscape [46]. However, they do not elucidate why the landscape is flat or how it relates to generalization and robustness, leaving the role of flatness in diffusion models unexplored. While prior works [1, 6, 32] leverage smoothness via Lipschitz continuity or gradient regularization, they do not directly examine flatness. To the best of our knowledge, we are the first to systematically investigate how flat minima influence diffusion models, their generative performance, and robustness–motivating our theoretical and empirical analysis.
6 Discussion and Limitation
The generative modeling community has long sought to understand generalization in terms of preventing models from overfitting to the training set while ensuring the generation of diverse and high-quality samples, particularly for underrepresented classes [35, 47]. While our work primarily investigates the role of flat minima in addressing exposure bias and quantization error,
FID $\downarrow$
ADM 15.97 (+5.34)
+SAM 11.72 (+4.72)
+SAM
@@ 6 Discussion and Limitation
7 Conclusion
We presented an in-depth investigation of the role and benefits of flat loss surfaces in generative models, which have remained unknown despite the great success of deep generative models. Our study explored this unknown by analyzing how flatness influences generative modeling and demonstrating its possible impact on robustness. First, our theoretical analysis reveals that the loss flatness of the generative models enhances the robustness against the disruptions of the target data distributions or the perturbations of model parameters. These insights naturally link to the reduced exposure bias problem and the model quantization error. Based on the evaluations on CIFAR-10, LSUN Tower, and FFHQ datasets, we demonstrated that 1) SAM is suitable to seek flatter minima of diffusion models rather than SWA and EMA, 2) flat minima reduce exposure bias with a minimal gap between testing and training, and 3) a flat model keeps its generative performance even with a strong model quantization; coinciding with our theory.
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.2022R1C1C100849612, No.RS-2024-00459023), Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT):(No.RS-2020-II201336, Artificial Intelligence Graduate School Program at UNIST), (RS-2025-25442824, AI Star Fellowship Program at UNIST), (No.2022-0-00959, No.RS-2022-II220959 (Part 2) Few-Shot Learning of Causal Inference in Vision and Language for Decision Making), (RS-2025-25441996, Development of a Virtual Tactile Signal Generation Platform Technology Based on Multimodal Vision-Tactile Integrated AGI), (RS-2025-25442149), Korea Health Technology R&D Project through the Korea Health Industry Development Institute (KHIDI), funded by the Ministry of Health & Welfare, Republic of Korea (No.RS-2025-02223382) and 2025 Research Fund (1.250006.01) of UNIST(Ulsan National Institute of Science & Technology.
Appendix
Appendix A Mathematical claims and proofs
@@
[…]
@@
[…]
@@
where $I(\mathbf{x},\delta):=\frac{1}{2}\mathbf{x}^{\top}(\delta\mathbf{W}^{\top})\mathbf{x}+\mathbf{x}^{\top}\delta(\mathbf{U}^{\top}\mathbf{e})+C$, and $C\in\mathbb{R}$ is set to satisfy $\int_{\mathbb{R}^{d}}e^{-I(\mathbf{x},\delta)}p(\mathbf{x})d\mathbf{x}=1$.
@@
$s_{\boldsymbol{\theta}}(\mathbf{x},t):=\frac{1}{m}\boldsymbol{\theta}\sigma(\mathbf{W}^{\top}\mathbf{x}+\mathbf{U}^{\top}\mathbf{e}_{t})$ (20)
@@
[…]
@@
[…]
@@
$=\frac{1}{m}(\boldsymbol{\theta}+\delta)\sigma(\mathbf{W}^{\top}\mathbf{x}+\mathbf{U}^{\top}\mathbf{e})-\nabla_{\mathbf{x}}\log{p(\mathbf{x})}$ (24)
$=\frac{1}{m}\boldsymbol{\theta}\sigma(\mathbf{W}^{\top}\mathbf{x}+\mathbf{U}^{\top}\mathbf{e})+\frac{1}{m}\delta\sigma(\mathbf{W}^{\top}\mathbf{x}+\mathbf{U}^{\top}\mathbf{e})-\nabla_{\mathbf{x}}\log{p(\mathbf{x})}$ (25)
@@
$\frac{1}{m}\delta(\mathbf{W}^{\top}\mathbf{x}+\mathbf{U}^{\top}\mathbf{e})-\nabla_{\mathbf{x}}\log{p(\mathbf{x})}$ (26)
@@
$\frac{\partial I(\mathbf{x})}{\partial\mathbf{x}}=\frac{1}{m}\delta(\mathbf{W}^{\top}\mathbf{x}+\mathbf{U}^{\top}\mathbf{e})$ (27)
@@
$I(\mathbf{x})=\frac{1}{2}\mathbf{x}^{\top}(\delta\mathbf{W}^{\top})\mathbf{x}+\mathbf{x}^{\top}\delta(\mathbf{U}^{\top}\mathbf{e})+C,$ (28)
with the assumption $\delta\mathbf{W}^{\top}$ is symmetric and where $C$ is a constant real number.
$\frac{1}{m}\delta(\mathbf{W}^{\top}\mathbf{x}+\mathbf{U}^{\top}\mathbf{e})-\nabla_{\mathbf{x}}\log{p(\mathbf{x})}$ (29)
@@
$=\exp{\Big{\{}-\frac{1}{2m}\mathbf{x}^{\top}(\delta\mathbf{W}^{\top})\mathbf{x}-\frac{1}{m}\mathbf{x}^{\top}\delta(\mathbf{U}^{\top}\mathbf{e})-C)\Big{\}}}p(\mathbf{x})$ (34)
[…]
@@
  • $\epsilon(\mathbf{x})$: the true Gaussian distribution that is known in the training process.
  • $\hat{\epsilon}(\mathbf{x})$: the perturbed distribution that caused by the $\delta$ model perturbation in Eq. (33).
@@
$\hat{\epsilon}(\mathbf{x})$ $:=e^{-I(\mathbf{x},\delta)}\epsilon$ (36)
$=\exp\bigg{(}-\frac{1}{2m}\mathbf{x}^{\top}\delta_{w}\mathbf{x}-\frac{1}{m}\mathbf{x}^{\top}\delta_{u}-C\bigg{)}\cdot\frac{1}{\sqrt{(2\pi)^{d}}}\exp\bigg{(}-\frac{1}{2}\mathbf{x}^{\top}\mathbf{x}\bigg{)},$ (37)
 $\qquad\text{ where }\mathbf{\delta}_{w}:=\delta\mathbf{W}^{\top},\delta_{u}:=\delta\mathbf{U}^{\top}\mathbf{e},\text{ and }\epsilon\sim\mathcal{N}(0,\mathbf{I})$
$=\frac{1}{\sqrt{(2\pi)^{d}}}\exp\bigg{(}-\frac{1}{2}\mathbf{x}^{\top}(\mathbf{I}+\frac{1}{m}\mathbf{\delta}_{w})\mathbf{x}-\frac{1}{m}\mathbf{x}^{\top}\delta_{u}-C\bigg{)}$ (38)
$=\frac{1}{\sqrt{(2\pi)^{d}}}\exp\bigg{(}-\frac{1}{2}\mathbf{x}^{\top}(\mathbf{I}+\frac{1}{m}\mathbf{\delta}_{w})\mathbf{x}-\frac{1}{m}\mathbf{x}^{\top}\delta_{u}-\frac{1}{2m^{2}}\delta_{u}^{\top}\Sigma_{\delta}\delta_{u}-\frac{1}{2}\log|\Sigma_{\delta}|\bigg{)},$ (39)
 $\qquad\text{ where }C=\frac{1}{2m^{2}}\delta_{u}\Sigma_{\delta}\delta_{u}+\frac{1}{2}\log|\Sigma_{\delta}|$ (40)
$=\frac{1}{\sqrt{(2\pi)^{d}|\Sigma_{\delta}|}}\exp\bigg{(}-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu}_{\delta})^{\top}{\Sigma_{\delta}}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{\delta})\bigg{)},\text{ where }{\Sigma_{\delta}}:=\bigg{(}\mathbf{I}+\frac{\delta_{w}}{m}\bigg{)}^{-1},\text{ and }\boldsymbol{\mu}_{\delta}:=-\frac{1}{m}\Sigma_{\delta}\delta_{u}$ (41)
Because the $\hat{\epsilon}(\mathbf{x})$ is also the Gaussian distribution, we present the KL Divergence between $\hat{\epsilon}(\mathbf{x})=\mathcal{N}(\boldsymbol{\mu}_{\delta},\Sigma_{\delta})$ and ${\epsilon}(\mathbf{x})=\mathcal{N}(0,\mathbf{I})$ as follows: ∎
@@
$\mathcal{E}$ $\leq\max_{\|\delta\|_{2}\leq\Delta}\frac{1}{2}\Bigg{[}\log|\Sigma_{\delta}|-d+tr(\Sigma_{\delta}^{-1})+{\boldsymbol{\mu}_{\delta}}^{\top}\Sigma_{\delta}^{-1}\boldsymbol{\mu}_{\delta}\Bigg{]}$ (46)
$\leq\frac{1}{2}\Bigg{[}\sum_{i}^{d}(\sigma_{i}-\log\sigma_{i})-d+\frac{\sigma_{d}}{m^{2}}\|\mathbf{U}^{\top}\mathbf{e}\|_{2}^{2}\Delta^{2}\Bigg{]}$ (47)
@@
[…]
$=\max_{\|\delta\|_{2}\leq\Delta}\frac{1}{2}\Bigg{[}\log|\Sigma_{\delta}|-d+tr(\Sigma_{\delta}^{-1})+{\boldsymbol{\mu}_{\delta}}^{\top}{\Sigma_{\delta}}^{-1}\boldsymbol{\mu}_{\delta}\Bigg{]}$ (49)
$=\max_{\|\delta\|_{2}\leq\Delta}\frac{1}{2}\Bigg{[}\sum_{i}^{d}\log\frac{1}{\sigma_{i}}-d+\sum_{i}^{d}\sigma_{i}+{\boldsymbol{\mu}_{\delta}}^{\top}{\Sigma_{\delta}}^{-1}\boldsymbol{\mu}_{\delta}\Bigg{]}$ (50)
$\leq\frac{1}{2}\Bigg{[}\sum_{i}^{d}(\sigma_{i}-\log\sigma_{i})-d+\frac{\sigma_{d}}{m^{2}}\|\mathbf{U}^{\top}\mathbf{e}\|_{2}^{2}\Delta^{2}\Bigg{]},$ (51)